OpenStudy (anonymous):

Solve for x. 6x - 15 = 9x - 6

sam (.sam.):

\[\begin{array}{l} 6 x-15=9 x-6 \\ \text{Add }(15-9 x)\text{ to both sides:} \\ -3 x=9 \\ \text{Divide both sides by }-3: \\ x=-3 \\\end{array}\]

OpenStudy (anonymous):

9x-6x=-15+6 ,3x=-9,x=-3

OpenStudy (anonymous):

oh yaa it is -3 i did not divide -9 by 3

OpenStudy (anonymous):

thankkk youuuuuuuuuuuuuuuuu